A simple pendulum with a solid bob has period T. what wil be the period of it if it is made to oscilate in a non viscous liquid of density §?(density of bob is d.)

Grade:12

1 Answers

vikas askiitian expert
13 years ago

T is inversly proportional to sqrt of weight

T  =  KW-1/2 

T/T1 = (W/W1)-1/2

W=mg = dVg                 (v is the volume of bob ,d dencity of bob & dw dencity of liquid)

W1 = dVg - dwVg

substituting these values

T1 = T[(d - dw)/d]-1/2

T1 = T / [1 - dw/d]1/2